A life jacket (personal flotation device) worn by commercial fishermen is designed to provide a minimum upward force on the user of 34 lbf. The foam material used in this jacket has a density of 3.2 lbm/ft3. If the remaining material (cloth, straps, fasteners, etc.) weighs 4.0 lbf and is of negligible volume, determine the minimum volume of the foam needed (ft3) for this life jacket.
Answer #1

This problem is from the fluid statics and we know that the water exerts a upward force on any object when it is partially orDensity of water is 62.4 lb ft^3. Thus m=X*624 lb F1 = 62.4x lbf F3=4 lbf F2 = 3.2*x Ibf In the limiting case when we need mi